Devon, Frank, Fua, Morris, and Pedro belong to the same gym. The gym rules prohibit the staff from giving out personal information about their clients. However, each of the five men said just enough that you can figure out their exact weights.
Devon: Pedro weighs 18 pounds more than I do. None of us weighs over 200 pounds.
Frank: My weight is divisible by 7. Morris weighs 12 pounds more than I do.
Fua: Three people are heavier than I am. Pedro's weight is a prime number.
Morris: I'm heavier than Pedro. Devon is the lightest.
Pedro: Morris's weight is divisible by 10. The five of us together weigh a total of exactly 840 pounds.
This is a problem from the textbook "Crossing the River with Dogs" but I'm having a difficult time answering using a problem solving strategy. Thanks!

Answer by richwmiller(17219)

Devon - 155
Fua - 164
Frank - 168 7*24=168
Pedro - 173 155+18 (18 more than Devon, and a prime number)
Morris - 180 168+12 180/10=18 (12 more than Frank, and divisible by 10)
Total of 840, all rules met.